### 20 Mar 2024 v1

# Decentralized Federated Learning: Model Update Tracking Under Imperfect Information Sharing

Abstract
A novel Decentralized Noisy Model Update Tracking Federated Learning algorithm (FedNMUT) is proposed, which is tailored to function efficiently in the presence of noisy communication channels that reflect imperfect information exchange. This algorithm uses gradient tracking to minimize the impact of data heterogeneity while minimizing communication overhead. The proposed algorithm incorporates noise into its parameters to mimic the conditions of noisy communication channels, thereby enabling consensus among clients through a communication graph topology in such challenging environments. FedNMUT prioritizes parameter sharing and noise incorporation to increase the resilience of decentralized learning systems against noisy communications. Through theoretical and empirical validation, it is demonstrated that the performance of FedNMUT is superior compared to the existing state-of-the-art methods and conventional parameter-mixing approaches in dealing with imperfect information sharing. This proves the capability of the proposed algorithm to counteract the negative effects of communication noise in a decentralized learning framework.
Index Terms:
Optimization algorithms, Distributed control, Stochastic systems, Numerical algorithms
I Introduction
In contemporary real-world scenarios, vast amounts of data are generated from various sources such as computers, mobile devices, smartwatches, and vehicles. These data are aggregated in centralized data centers for the purpose of training machine learning models. However, this centralized approach encounters significant hurdles, including limited communication bandwidth and privacy concerns, rendering it unreliable and non-scalable. Consequently, there is a pressing need to establish learning paradigms that not only ensure data security but also uphold privacy standards. This necessity has spurred the development of decentralized optimization algorithms, exemplified by Decentralized Federated Learning (DFL) and Federated Learning (FL). These methodologies have been increasingly applied across diverse sectors, including smart cities and connected vehicles [1, 2, 3, 4, 5, 6, 7, 8, 9].
I-A Literature Overview
Decentralized optimization typically employs consensus-based gradient descent methods, where the computed parameters are shared among clients [10, 1, 11]. These clients independently calculate local weights and gradients based on their data, subsequently exchanging these parameters with others. The aggregation of these parameters, influenced by the network’s topology which governs the communication framework, is pivotal in this learning paradigm. This network topology is often depicted as a simple graph, illustrating the communication pathways among clients. Decentralized approaches aim to alleviate the limitations inherent in centralized systems, such as communication delays and bandwidth constraints, thereby enhancing scalability and efficiency in extensive networks.
Recent advancements like Gradient Tracking (GT) and Momentum Tracking (MT) have been devised to tackle the challenge of data heterogeneity in decentralized settings, albeit at the cost of increased communication overhead [12, 13, 14, 15]. Conversely, the Quasi-Global Momentum (QGM) strategy achieves global momentum synchronization without necessitating additional communication, thus mitigating decentralized learning challenges related to diverse data sets. Furthermore, RelaySGD, a novel method, replaces traditional gossip averaging with Relay-Sum, offering a unique approach to the averaging process. The integration of RelaySGD with existing methods can significantly improve performance. The QG-DSGDm algorithm, which combines QGM with DSGDm, marks a notable advancement in this domain. A recent proposition, the Global Update Tracking (GUT) method, addresses data heterogeneity in decentralized learning efficiently, without incurring extra communication costs. Empirical results validate that GUT not only accommodates variances in data distribution across devices but also bolsters the performance of decentralized learning processes [16].
While enhancing communication efficiency remains a crucial challenge in distributed learning, efforts like communication compression have been explored [17, 18, 3, 19, 20, 21]. However, these initiatives often presume ideal, noise-free communication channels. The robustness and dependability of machine learning frameworks, especially in the burgeoning fields reliant on distributed learning, hinge on their performance amidst noisy communications.
Imperfect information exchange, such as noisy or quantized communication, has been examined in the context of average consensus algorithms within distributed frameworks. Yet, the ramifications of varying noise levels remain underexplored. Moreover, existing research, primarily focused on consensus issues, does not fully address the complex challenges encountered in contemporary decentralized optimization and learning paradigms [22, 23]. In contrast to Federated Learning (FL), where server assistance is common, Decentralized Federated Learning (DFL) operates without a central server, with each client acting autonomously, processing local Stochastic Gradient Descent (SGD) or its variants on its data and interacting directly with neighboring clients.
Our preceding work delineated three algorithms for DFL under imperfect communication conditions, typified by noisy channels. These algorithms—FedNDL1, FedNDL2, and FedNDL3—differ in their handling of noise and parameter sharing, showcasing varying degrees of resilience to communication noise [24].
This paper introduces a novel algorithm that employs the Gradient Tracking method in DFL, considering the impact of communication noise. Previous studies have assessed the efficacy of two-time scale methods in DFL with noisy channels, but these investigations were constrained by rigid assumptions that are seldom met in practical, large-scale learning scenarios [25, 26, 27, 28, 29].
I-B Paper’s Contributions
We evaluate the performance of decentralized FL using the Federated Model Update Tracking algorithm, which incorporates noise into tracking parameter transmissions to clients or servers. Unlike prior models where communication noise was not considered, our approach integrates noise post-local SGD updates, facilitating parameter exchange via a gossip/mixing matrix and updating the global parameters thereafter. We also compare the noise resilience of our algorithm with previously developed algorithms, providing both theoretical and empirical evidence of its robustness against communication noise [24].We show that the proposed algorithm handles the communication noise well than the FedNDL3 algorithm.
II Problem Statement
In this section, we describe the problem structure, assumptions, and the proposed algorithms that we analyzed in this paper. We start with a standard DFL setup in which $n$ clients have their own local datasets and collaborate with each other to update the global parameters. Formally, the problem can be represented as
$\min_{x\in\mathbb{R}^{d}}\Big{[}f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x_{i})\Big{]},$ (1)
where $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ for $i\in\{1,\dots,n\}$ is the local objective function of the $i^{th}$ client node. The stochastic formulation of the local objective function can be written
$f_{i}(x_{i})=\operatorname{\mathbb{E}}_{\xi_{i}\sim\mathcal{D}_{i}}[\ell(x_{i},\xi_{i})],$ (2)
where $\xi_{i}$ is the data that has been sampled from the data distribution $\mathcal{D}_{i}$ for the $i^{th}$ client. The function $\ell(x_{i},\xi_{i})$ is the loss function evaluated for each client and for each data sample $\xi_{i}$. Here $x_{i}\in\mathbb{R}^{d}$ is the parameter vector of client $i$, and $X\in\mathbb{R}^{d\times n}$ is the matrix formed using these parameter vectors. The primary objective of the clients is to achieve optimality through collaboration i.e., $x_{i}=x^{*}=\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}f(x)$, which is a global minimizer.
Definition 1 (Mixing matrix).
The mixing/gossip matrix, $W=[w_{ij}]\in[0,1]^{n\times n}$, is a non-negative, symmetric $(W=W^{\top})$ and doubly stochastic $(W\mathds{1}=\mathds{1},\mathds{1}^{\top}W=\mathds{1}^{\top})$ matrix, where $\mathds{1}$ is the column vector of unit elements of size $n$
FedNDL1 [24]: In this algorithm, each client in parallel performs updates first and then communicates the updated parameters to their neighbors. The communication depends on the topology of the communication graph, i.e., the mixing matrix, $W$, through a noisy communication channel. Due to the noisy communication channel, the neighboring client receives a noisy version of the parameters,
$x_{i}^{(t+1)}\ =\ \sum_{j=1}^{n}w_{ij}\ (x_{j}^{(t+\frac{1}{2})}+\delta_{j}^{(t)}),$ (3)
where $\delta_{j}^{(t)}\in\mathbb{R}^{d}$, is a zero mean random noise and $x_{j}^{(t+\frac{1}{2})}$ is the vector of parameters sent by client $j$. Since we assume the noise to have a zero mean, the noise variance is
$D^{2}_{t,j}=\operatorname{\mathbb{E}}[\|\delta_{j}^{(t)}\|^{2}].$ (4)
FedNDL3 [24]: In this algorithm, the clients share their gradients over a noisy communication channel instead of the weights followed by the SGD update. This idea comes from our Noisy-FL motivation and the fact that SGD is inherently a noisy process. So, pursuing this scenario gives more flexibility to handle the noise as a part of the SGD process.
$x_{i}^{(t+1)}\ =\ x_{i}^{(t)}-\eta_{t}\sum_{j=1}^{n}w_{ij}\ (g_{j}^{(t)}+\delta_{j}^{(t)}),$ (5)
where $g_{j}^{(t)}$ refers to the gradient of client $j$ at iteration $t$
FedNMUT: We next describe our proposed algorithm studied in this paper.
$x_{i}^{(t+1)}=x_{i}^{(t)}-\eta_{t}\widetilde{y_{i}}^{t}\$
$x_{i}^{(t+1)}=x_{i}^{(t)}-\eta_{t}({y_{i}}^{t}+\delta_{i}^{t})\$ (6)
$y_{i}^{t}=\Delta_{i}^{t}+\mu\Big{[}\sum\limits_{j\in\mathcal{N}(i)}w_{ij}(\widetilde{y_{j}}^{t-1}-\frac{1}{\eta}(\hat{x}_{j}^{t}-x_{i}^{t}))-\Delta_{i}^{t-1}\Big{]}\$
$\Delta_{i}^{t}=g_{i}^{t}-\frac{1}{\eta}\sum_{j\in\mathcal{N}(i)}w_{ij}(\hat{x}_{j}^{t}-x_{i}^{t})$
In this algorithm, $\mu$ is FedNMUT scaling factor/hyper-parameter that needs to be tuned to provide optimal solution. As noticed in Figure  1 and Figure 2, the convergence of the algorithm is dependent on the value of $\mu$.
Algorithm FedNMUT - Noisy Model Update Tracking
1:  Input: For each node $i$ initialize: $x_{i}^{(0)}\ \in\ \mathbb{R}^{d}$, step size $\{\eta_{t}\}_{t=0}^{T-1}$, neighbors’ copy $\hat{x}_{j}^{0}$, step size $\eta$, GUT scaling factor $\mu$, mixing matrix $W=[w_{ij}]_{i,j\in[1,n]}$, $\mathcal{N}(i)$ represents neighbors of $i$ including itself, noise from the communication channel $\delta^{(t)}$
2:  for $t=0,\dots,T-1$ do
3:      Run in parallel for each client $i$
4:      Sample $\xi_{i}^{(t)}\mbox{, compute }g_{i}^{(t)}=\widetilde{\nabla}f_{i}(x_{i}^{(t)},\xi_{i}^{(t)})$
5:      $\Delta_{i}^{t}=g_{i}^{t}-\frac{1}{\eta}\sum_{j\in\mathcal{N}(i)}w_{ij}(\hat{x}_{j}^{t}-x_{i}^{t})$
6:     $y_{i}^{t}=\Delta_{i}^{t}+\mu\Big{[}\sum\limits_{j\in\mathcal{N}(i)}w_{ij}(\widetilde{y_{j}}^{t-1}-\frac{1}{\eta}(\hat{x}_{j}^{t}-x_{i}^{t}))-\Delta_{i}^{t-1}\Big{]}$
7:     $\widetilde{y_{i}}^{t}=y_{i}^{t}+\delta_{i}^{(t)}$
8:     SENDRECEIVE($\widetilde{y_{i}}^{t}$)
9:     $x_{i}^{t+1}=x_{i}^{t}-\eta\widetilde{y_{i}}^{t}$
10:     $\hat{x}_{j}^{t+1}=\hat{x}_{j}^{t}-\eta\widetilde{y_{j}}^{t}\hskip 5.69054pt\forall\hskip 5.69054ptj\in\mathcal{N}(i)\backslash i$
11:  end for
(a) Fully-connected topology
(b) Torus topology
(c) Ring topology
Figure 1: Loss vs. iterations without noise for various $\mu$ values for different communication topologies.
(a) Fully-connected topology
(b) Torus topology
(c) Ring topology
Figure 2: Consensus error vs. iterations without noise for various $\mu$ values for different communication topologies.
II-A Assumptions
We now discuss the assumptions made in our analysis of the algorithms. They are standard assumptions used in the analysis of decentralized algorithms, see [30, 2, 3].
Assumption 1 (Smoothness).
The objective function $\ell(x,\xi)$ is $L$-smooth with respect to $x$, for all $\xi$. Each $f_{i}(x)$ is $L$-smooth, that is,
[…]
Hence the function $f$ is also $L$-smooth.
Assumption 2 (Bounded Variance).
The variance of the stochastic gradient of each client $i$ is bounded,
$\operatorname{\mathbb{E}}[||\widetilde{\nabla}f_{i}(x_{i}^{t},\xi_{i}^{t})-\nabla f_{i}(x_{i}^{t})||^{2}]\leq\sigma^{2},$
where $\xi_{i}^{t}$ denotes random batch of samples in client node $i$ for $t^{th}$ round, and ~$\widetilde{\nabla}f_{i}(x_{i}^{t},\xi_{i}^{t})$ denotes the stochastic gradient. In addition, we also assume that the stochastic gradient is unbiased, i.e., $\operatorname{\mathbb{E}}[\widetilde{\nabla}(f_{i}(x_{i}^{t},\xi_{i}^{t}))]=\nabla f_{i}(x_{i}^{t})$.
Assumption 3 (Mixing matrix).
The mixing matrix W satisfies for $\rho\in(0,1]$,
$\|(\bar{X}-X)W\|_{F}^{2}\leq(1-\rho)\|\bar{X}-X\|_{F}^{2},$
which means that the gossip averaging step brings the columns of $X\in\mathbb{R}^{d\times n}$ closer to the row-wise average, that is, $\bar{X}=X\frac{\mathds{1}\mathds{1}^{\top}}{n}$.
Note that standard topologies such as ring, torus, and fully-connected satisfy the above assumption.
Assumption 4 (Bounded Client Dissimilarity (BCD)).
For all $x\in\mathbb{R}^{d}$, where $\zeta$ is a constant.
$\frac{1}{n}\sum_{i=1}^{n}\|\nabla f_{i}(x)-\nabla f(x)\|^{2}\leq\zeta^{2},$
The above assumption is made to limit the extent of client heterogeneity and is standard in the DFL setup.
Assumption 5 (Noise model).
The noise present due to contamination of communication channel $\delta_{i}^{(t)}$ is independent, has zero mean and bounded variance, that is, $\operatorname{\mathbb{E}}[\delta_{i}^{(t)}]=0$ and $\operatorname{\mathbb{E}}[||\delta_{i}^{(t)}||^{2}]=D_{i}^{2,t}<\infty$.
This assumption is specific to the imperfect information sharing setup and is considered recently in [27, 28, 31].
(a) Fully-connected topology
(b) Torus topology
(c) Ring topology
Figure 3: Loss vs. iterations with and without noise (Var=0.005) for different communication topologies. $\mu$ = 0.02
(a) Fully-connected topology
(b) Torus topology
(c) Ring topology
Figure 4: Consensus error vs. iterations with and without noise (Var=0.005) for different communication topologies. $\mu$ = 0.02. Note the different Y-axis scale in Fig (a) as compared with Fig (b) and (c) for better readability.
III Convergence Analysis
In this section, we state the main theorem providing an upper bound on the convergence errors of FedNMUT. The convergence results are for non-convex $L$-smooth loss functions and noisy communication channels.
Following are the major lemmas derived during the convergence analysis of the proposed algorithm. Detailed proof and derviation of the lemmas and theorem are provided in the full journal version of the paper.
Lemma 1.
Given assumptions 1-3 and $\eta\leq\frac{1}{4L}$, we have
$\mathbb{E}f(\bar{x}^{t+1})\leq\mathbb{E}f(\bar{x}^{t})+\frac{L\eta^{2}\sigma^{2}}{2n}-\frac{3\eta}{8}\mathbb{E}||\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(x_{i}^{t})||^{2}$
$-\frac{\eta}{2}\mathbb{E}||\nabla f(\bar{x}^{t})||^{2}+\frac{L^{2}\eta}{2n}\mathbb{E}||X^{t}-\bar{X}^{t}||_{F}^{2}+\frac{L\eta^{2}}{2n}\sum_{i=1}^{n}D^{2}_{t,i}$
Lemma 2.
Given assumptions 1-3 and $\eta\leq\frac{\rho}{7L}$, we have
$\frac{1}{n}\mathbb{E}||X^{t+1}-\bar{X}^{t+1}||_{F}^{2}\leq\frac{1-\rho/4}{n}\mathbb{E}||X^{t}-\bar{X}^{t}||_{F}^{2}+\frac{4\eta^{2}\zeta^{2}}{\rho}$
$+\frac{4\eta^{2}\sigma^{2}}{n}+\frac{6\eta^{2}\mu^{2}}{\rho n}\mathbb{E}||B^{t}||_{F}^{2}+\frac{4\eta^{2}}{n\rho}D^{2,t}$
Lemma 3.
Given assumptions 1-3 and $\frac{\mu}{1-\mu}\leq\frac{\rho}{42}$, we have
$\frac{6\eta^{2}\mu^{2}}{\rho n(1-\mu)}\mathbb{E}||B^{t+1}||_{F}^{2}\leq\Big{(}\frac{6\eta^{2}\mu^{2}}{n\rho(1-\mu)}-\frac{6\eta^{2}\mu^{2}}{n\rho}\Big{)}\mathbb{E}||B^{t}||_{F}^{2}$
$\hskip 11.38109pt+\frac{\rho}{8n}\mathbb{E}||X^{t}-\bar{X}^{t}||_{F}^{2}+\frac{\eta^{2}\rho\sigma^{2}(1-\mu)}{8}+\frac{\eta^{2}\rho D^{t,2}}{8}+\frac{\eta^{2}\rho\zeta^{2}}{8}$
Theorem 1 (Smooth non-convex cases for FedNMUT).
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}||\nabla f(\bar{x}^{t})||^{2}\leq\frac{2}{\eta T}(f(\bar{x}^{0}-f^{*})$
$\hskip 7.11317pt+2L^{2}\eta^{2}\sigma^{2}\Big{[}\frac{16}{\rho n}+\frac{1-\mu}{2}+\frac{1}{2nL\eta}\Big{]}+2L^{2}\eta^{2}\zeta^{2}\Big{[}\frac{48}{\rho^{2}}+\frac{1}{2}\Big{]}$
$\hskip 7.11317pt+2L^{2}\eta^{2}D^{t,2}\Big{[}\frac{16}{n\rho^{2}}+\frac{1}{2}+\frac{1}{2nL\eta}\Big{]}$
where all expectations are w.r.t. the data and the noise.
Corollary 1.1.
Suppose that the step size satisfies $\eta=\mathcal{O}\Big{(}\sqrt{\frac{n}{T}}\Big{)}$ For a sufficiently large $T$ we have,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}||\nabla f(\bar{x}^{t})||^{2}\leq\mathcal{O}\Bigg{(}\frac{1}{\sqrt{nT}}+\frac{1}{T}\Bigg{)}.$ (8)
Theorem 1 establishes a worst-case upper bound on the convergence of the algorithm studied in the paper. In particular, the theorem jointly bounds the expected gradient norm, which is a notion of approximate first-order stationarity of the average iterate $\bar{x}_{t}$, and the consensus error. The convergence bounds consist of four terms: the first term effectively captures the error arising from inaccurate initialization. The second and third term captures the effect of stochasticitiy and data heterogeneity. The last term captures the adverse effect of imperfect communication modeled as communication noise.
The finding presented in Corollary 1.1 shows that the FedNMUT algorithm can achieve a linear speedup, exhibiting a convergence rate of $\mathcal{O}(\frac{1}{\sqrt{nT}})$, given that $T$ is large enough and unaffected by the communication topology. This rate of convergence is comparable to the best-known results for decentralized SGD algorithms found in existing literature.
(a) Fully-connected topology
(b) Torus topology
(c) Ring topology
Figure 5: Loss vs. iterations with and without noise (Var=0.01) for different communication topologies. $\mu$ = 0.02
(a) Fully-connected topology
(b) Torus topology
(c) Ring topology
Figure 6: Consensus error vs. iterations with and without noise (Var=0.01) for different communication topologies. $\mu$ = 0.02. Note the different Y-axis scale in Fig (a) as compared with Fig (b) and (c) for better readability.
IV Experiments
We conduct a series of experiments on regression tasks to evaluate the effects of noise on the convergence behavior of the algorithm outlined in Theorem 1. We set the number of clients, denoted as $n$, to 16. Each experiment is executed thrice, with the outcomes (loss/consensus error) being averaged. Our approach employs a mean-squared error loss function enhanced with $L_{2}$ regularization. We establish the initial learning rate at 0.2, applying a decay factor of 0.9 in subsequent iterations. We create data samples ($m$ = 10,000) in the form of ${(x_{i};y_{i})}^{m}_{i=1}$, modeled as $y_{i}=\langle w,x_{i}\rangle+\epsilon_{i}$, where $w$ belongs to $\mathbb{R}^{2000}$, $x_{i}$ follows a normal distribution $\mathcal{N}(0;I_{2000})$, and noise $\epsilon_{i}$ adheres to $\mathcal{N}(0,0.05)$.
Our experimentation spans various noise variances, $D_{t,i}^{2}$, across all $t,i$, reflecting the conditions specified in the algorithm. These tests encompass different communication topologies, specifically ring, torus, and fully connected networks. In these topologies, the nonzero elements of the mixing matrix hold the values of $\frac{1}{3}$, $\frac{1}{5}$, and $\frac{1}{n}$, respectively.
We commence with noise-free trials to establish a baseline, subsequently incrementing noise variance to examine algorithmic robustness. For consistency, the experimental findings with noise variances of $D^{2,t}=0.005$ and $D^{2,t}=0.01$ are depicted alongside the no-noise condition in Figures 3, 4, 5 and 6. These tests were conducted using an Intel Xeon Gold workstation.
The consensus error depends on the topology of the communication network. We observed that the consensus error is low for the fully connected network and high for the ring topology for the same algorithm in the presence of noise which is also consistent with the Theorem 1.
V Conclusion
We studied the impact of noisy communication channels on the convergence of Decentralized Federated Learning with Model Update Tracking method. We proposed multiple scenarios for establishing consensus in the presence of noise and provided experimental results on the algorithm. Additionally, we provided theoretical results for FedNMUT under the assumption of smooth non-convex function, and we observed that the noise term in the upper bound given by Theorem 1 is of order $\mathcal{O}(\frac{1}{\sqrt{nT}})$ and is independent of communication topology. We conducted numerical experiments and observed that FedNMUT is more robust against the added noise than the existing state-of-the-art methods that include communication noise.